Determine the charge on the object and explain why is has this charge.

Object
1. An atom that has 16 electrons and 15 protons
2. An atoms that has 100 electrons and 120 protons
3. A neutral object that loses 100 electrons to the ground.
1
Expert's answer
2021-03-29T09:03:15-0400

The charge of one electron is


"e \approx -1.6\times 10^{-19}C"

The charge of one proton is


"p \approx 1.6\times 10^{-19}C"

1. 15 protons and 15 electrons compensate each other, and the total charge is caused by one extra electron. Thus:


"q_1 = e = -1.6\times 10^{-19}C"

2. 100 protons and 100 electrons compensate each other, and the total charge is caused by 20 extra protons:


"q_2 = 20p = 20\cdot 1.6\times 10^{-19}C = 3.2\times10^{-18}C"

3. Since the initially neutral object looses 100 electrons, it will have 100 not compensated protons. Thus, its total charge will be:


"q_3 = 100p = 100\cdot 1.6\times 10^{-19}C = 1.6\times10^{-17}C"

Answer. 1. "-1.6\times 10^{-19}C"

2. "3.2\times10^{-18}C"

3. "1.6\times10^{-17}C"